OpenStudy (anonymous):

f(x)=x^2+6x-16 g(x)=x^2+6x+1 two real irrational solutions two real rational solutions

OpenStudy (lunymoony):

Are you asking which one is a rational or irrational solution of the two given problems?

OpenStudy (anonymous):

f(x) (x +8)(x-2) two reals -8,+2 g(x) \[\frac{ -b \pm \sqrt{b ^{2}-4ac} }{ 2a }\]

OpenStudy (anonymous):

\[\frac{ -6\pm \sqrt{6^{2}-4*1*1} }{ 2*1 }=\frac{ -6\pm \sqrt{36-4} }{ 2}\] \[-3\pm \frac{ \sqrt{32} }{ 2 }=-3\pm2\sqrt{2}\]